OpenStudy (anonymous):

find the given derivative by finding the first few derivatives and observing the pattern that occurs : D^35 xsin pi x

OpenStudy (solomonzelman):

I am not sure what \(\large\color{black}{ \displaystyle {\rm D}^{35} }\) is, does that come to denote 35th derivative?

OpenStudy (solomonzelman):

In any case, \(\large\color{black}{ \displaystyle f(x)=x\sin\left(\pi x\right) }\) \(\large\color{black}{ \displaystyle f'(x)=\sin\left(\pi x\right)+\pi x\cos\left(\pi x\right) }\) \(\color{black}{ \displaystyle f''(x)=\pi\cos\left(\pi x\right)+\pi \cos\left(\pi x\right)-\pi^2 x\sin\left(\pi x\right) = 2\pi \cos\left(\pi x\right)-\pi^2 x\sin\left(\pi x\right) }\)

OpenStudy (anonymous):

Yes I think that it denotes the 35th derivative, I understand since the derivative resets every 4th derivative, the answer would be the 3rd derivative of the equation. (since 4 goes into 32 evenly)

OpenStudy (solomonzelman):

with the x in front of the sine, the derivatives don't reset.

OpenStudy (anonymous):

Oh So in this case it would just be the second derivative ?

OpenStudy (solomonzelman):

What do you mean?

OpenStudy (anonymous):

Because you said that it doesn't reset if there is an x in front

OpenStudy (solomonzelman):

well, not reset to the very first function, but I think the pattern is present

OpenStudy (solomonzelman):

I think I have got the pattern.

OpenStudy (anonymous):

How do you determine the pattern ?

OpenStudy (solomonzelman):

I was just taking derivatives and simplifying them.... it might be a pain in the neck, but this is how you do it.

OpenStudy (solomonzelman):

Are my notations making sense?

OpenStudy (anonymous):

I understand what you just did, but would that change the final answer for the original questions I asked?

OpenStudy (solomonzelman):

you have to observe the pattern for the derivatives. Every 4th derivative... I would observe every pattern for (4k)th (4k+1)th (4k+2)th (4k+3)th (and (4k+4)th is again restarting the cycle of (4k)th )

OpenStudy (anonymous):

35(sin(pix) + pi x cos (pi x)

OpenStudy (solomonzelman):

you need 35th, or that would be, (4k+3)th derivative. and percisely: (4×8+3)th

OpenStudy (anonymous):

which would make it the 3rd derivative of the equation ?

OpenStudy (solomonzelman):

this is what I got: My function is: \(\large\color{black}{ \displaystyle f^{(0)}(x)=x\sin\left(\pi x\right) }\) \(\large\color{blue}{ \displaystyle f^{(4k)}(x)=\pi^{4k-1}\left(\pi x \sin \pi x~-4k\cos\pi x\right) }\) \(\large\color{blue}{ \displaystyle f^{(4k+1)}(x)=\pi^{4k}\left((4k+1) \sin \pi x+\pi x \cos\pi x\right) }\) \(\large\color{blue}{ \displaystyle f^{(4k+2)}(x)=\pi^{4k+1}\left((4k+2) \cos \pi x-\pi x \sin \pi x\right) }\) \(\large\color{blue}{ \displaystyle f^{(4k+3)}(x)=-\pi^{4k+2}\left((4k+3) \sin \pi x+\pi x \cos\pi x\right) }\)

OpenStudy (solomonzelman):

these are the patterns I got.... took a while to type and check them, apologize.

OpenStudy (solomonzelman):

Try finding 1st, 2nd, 3rd, 4th, and so forth derivative.... (I would recommend factoring like I did), you will notice the pattern repeats every 4th time (for 4k, 4k+1, 4k+2, and 4k+3)

OpenStudy (anonymous):

Thank you so much for your help !!

OpenStudy (solomonzelman):

Patterns do take pain, but all you need is work... you can do it:) you welcome